### v1

# On the Regularization of Learnable Embeddings for Time Series Processing

Abstract
In processing multiple time series, accounting for the individual features of each sequence can be challenging. To address this, modern deep learning methods for time series analysis combine a shared (global) model with local layers, specific to each time series, often implemented as learnable embeddings. Ideally, these local embeddings should encode meaningful representations of the unique dynamics of each sequence. However, when these are learned end-to-end as parameters of a forecasting model, they may end up acting as mere sequence identifiers. Shared processing blocks may then become reliant on such identifiers, limiting their transferability to new contexts. In this paper, we address this issue by investigating methods to regularize the learning of local learnable embeddings for time series processing. Specifically, we perform the first extensive empirical study on the subject and show how such regularizations consistently improve performance in widely adopted architectures. Furthermore, we show that methods preventing the co-adaptation of local and global parameters are particularly effective in this context. This hypothesis is validated by comparing several methods preventing the downstream models from relying on sequence identifiers, going as far as completely resetting the embeddings during training. The obtained results provide an important contribution to understanding the interplay between learnable local parameters and shared processing layers: a key challenge in modern time series processing models and a step toward developing effective foundation models for time series.
1 Introduction
Collections of related time series characterize many applications of learning systems in the real world, such as traffic monitoring (Li et al., 2018; Yu et al., 2018), energy analytics (Dimoulkas et al., 2019; Gasparin et al., 2022), climate modeling (Ma et al., 2023; Chen et al., 2023), and biomedical data processing (Jarrett et al., 2021; Zhang et al., 2022). The success of deep learning in the associated tasks, e.g., forecasting (Shih et al., 2019; Benidis et al., 2022), imputation (Cao et al., 2018; Cini et al., 2022) and virtual sensing (Wu et al., 2021), relies on effectively modeling shared patterns across time series while also accounting for their individual characteristics (Benidis et al., 2022). In this context, models must rely on some attributes or positional encodings to tailor the processing to the target time series, with the risk of requiring long observation windows and high model capacity when those are not available (Salinas et al., 2020; Montero-Manso & Hyndman, 2021). Indeed, prior and positional information is often unavailable or insufficient to provide effective specialization. Notably, the study of methods to tune a shared time series model on a specific task is a major concern for the development of foundation models, i.e., backbone models trained on large collections of time series and then applied to specific target applications (Garza & Mergenthaler-Canseco, 2023; Liang et al., 2024). This problem is also particularly relevant in spatiotemporal forecasting when dealing with multiple time series coming from a sensor network (Bai et al., 2020; Cini et al., 2023b).
Among different methods, research has addressed the problem by looking into hybrid global-local architectures, i.e., architectures that combine global models with local learnable components (e.g., layers) specific to a target time series (Wang et al., 2019; Smyl, 2020). As a prominent example, Smyl (2020) won the M4 competition (Makridakis et al., 2020) by combining a shared recurrent neural network (RNN) with local exponential smoothing models fitted on each time series. In the context of spatiotemporal data, research has shifted towards the adoption of learnable embeddings (Bai et al., 2020; Shao et al., 2022), i.e., vectors of learnable parameters, to reduce the cost of learning (more complex) local processing blocks (Cini et al., 2023b). Each embedding, associated with a specific time series, is fed into the shared modeling architecture and trained end-to-end alongside it. These representations can go beyond simply encoding coordinates, as they can account for the dynamics of each time series w.r.t. the other sequences in the collection (Cini et al., 2023b). Parameters of this kind are analogous to the word embeddings used in natural language processing (NLP) (Mikolov et al., 2013; Peng et al., 2015). We will refer to them as local embeddings.
While local embeddings offer significant advantages, their adoption introduces potential drawbacks. In the first place, differently from NLP applications that operate on fixed size dictionary, the target time series might change over time, and new sequences might be added to the collection. Secondly, since embeddings are learned jointly with the entire architecture, co-adaptation with the shared layers might result in embeddings being used as simple sequence identifiers (Geirhos et al., 2020). This interdependence can hinder the flexibility of the shared processing blocks and limit their transferability. Existing works (Yin et al., 2022; Cini et al., 2023b; Prabowo et al., 2024) show evidence that constraining the structure of the embedding space can lead to transferability improvements. However, no prior work has systematically addressed and evaluated regularization methods for local embeddings in forecasting architectures.
To fill this void, we investigate the impact of regularizing the learning of local embeddings for time series within a selection of commonly used deep learning architectures. In particular, we perform an extensive empirical study comparing a variety of regularization strategies, ranging from standard approaches, such as weight penalties and dropout, to more advanced methods, e.g., clustering and variational regularization. Our analysis considers a range of scenarios, including transductive settings, transfer learning, and sensitivity analyses. Empirical results confirm that methods that prevent the co-adaptation of the local and global blocks are consistently among the best-performing approaches. To further validate this observation, we include in the analysis a forget-and-relearn strategy (Zhou et al., 2021) that periodically resets embeddings during training. Results w.r.t. such baseline further show that decoupling the learning of local and global parameters offers a good design principle for regularization strategies in this context. Our main findings can be summarized as follows.
  1. (F1)
    In both transductive and transfer learning settings, regularizing the local embeddings can provide clear performance improvements across several architectures.
  2. (F2)
    Regularization strategies that prevent the co-adaptation between global and local parameters consistently lead to larger performance gains.
  3. (F3)
    Finding (F2) can be used as a design principle for designing new regularization strategies to mitigate overfitting and improve transferability.
The regularization of local learnable embeddings emerges as an often neglected but central aspect, which, with negligible computational overhead, can lead to significant performance improvements across all the considered benchmarks. This makes our study an important missing piece for guiding practitioners and researchers in building and designing modern neural architectures for time series processing.
2 Preliminaries
This section introduces the notation and formalizes the problem of time series forecasting, with a focus on the class of deep learning architectures we consider in this paper.
2.1 Problem settings
Consider a collection ${\mathcal{D}}$ of $N$ time series, where each $i$-th time series consists of a sequence of $T$ observations $\{{\bm{x}}^{i}_{t}\in{\mathbb{R}}^{d_{x}}\}_{t=1}^{T}$. Specifically, we indicate as related time series a set of homogenous time series coming from the same domain but generated by different sources (e.g., different sensors). Examples include sales figures for different products, or energy consumption of various users. Time series might be acquired both asynchronously and/or synchronously; in the latter case, we denote the stacked $N$ observations at time step $t$ by the matrix ${\bm{X}}_{t}\in{\mathbb{R}}^{N\times d_{x}}$ and use the shorthands ${\bm{X}}_{t:t+T}$ to indicate the sequence of observations within the time interval $[t,t+T)$ and ${\bm{X}}_{<t}$ to indicate those up to time $t$, excluding $t$. Eventual (exogenous) covariates associated with each time series are denoted by ${\bm{u}}^{i}_{t}\in{\mathbb{R}}^{d_{u}}$ (${\bm{U}}_{t}\in{\mathbb{R}}^{N\times d_{u}}$). We focus on multistep-ahead time series forecasting, i.e., the problem of predicting the next $H$ future values for each $i$-th time series ${\bm{x}}^{i}_{t:t+H}$ given exogenous variables and a window of $W$ past observations. We focus on point forecasts, while probabilistic predictors might also be considered.
2.2 Hybrid global-local architectures for time series
Figure 1: Overview of the hybrid global-local time series forecasting framework.
We consider a broad class of models similar to those in (Benidis et al., 2022) and (Cini et al., 2023a). In particular, we consider predictors such as
$\hat{\bm{x}}^{i}_{t:t+H}=F({\bm{x}}^{i}_{t-W:t},{\bm{u}}^{i}_{t-W:t+H};\theta),\quad i=1,...,N$ (1)
where $\theta$ are the learnable parameters of the model, $\hat{\bm{x}}^{i}_{t:t+H}$ indicates the predicted values of ${\bm{x}}^{i}_{t:t+H}$ and $F(\cdot;\theta)$ is a family of parametric forecasting models. Models in Eq. 1 do not take into account (spatial) dependencies that might exist among time series. In scenarios such as spatiotemporal forecasting, where such dependencies might be relevant to achieve accurate predictions, models that can operate on sets of (synchronous) correlated time series should be preferred. In particular, we consider a family of models
$\widehat{\bm{X}}^{{\mathcal{S}}}_{t:t+H}=F({\bm{X}}^{{\mathcal{S}}}_{t-W:t},\mathbf{U}^{{\mathcal{S}}}_{t-W:t+H};\theta),\quad\quad\forall{\mathcal{S}}\subseteq{\mathcal{D}}$ (2)
where $F(\cdot;\theta)$ operates on subsets ${\mathcal{S}}$ of the collection of time series ${\mathcal{D}}$ and ${\bm{X}}_{t}^{{\mathcal{S}}}\in{\mathbb{R}}^{|{\mathcal{S}}|\times d_{x}}$ the resulting stack of observations at time step $t$. Models of this kind can be implemented by architectures operating on sets (Zaheer et al., 2017), e.g., attention-based architectures (Vaswani et al., 2017; Grigsby et al., 2021) or spatiotemporal graph neural networks (STGNNs) (Jin et al., 2023; Cini et al., 2023a) based on message passing (Gilmer et al., 2017). These models can eventually account for priors on existing dependencies among time series, which could be encoded, e.g., by a graph adjacency matrix $\mathbf{A}\in{\mathbb{R}}^{N\times N}$ (Cini et al., 2023a). Note that models in both families (Eqs. 1 and 2) are global, i.e., they share parameters among all the time series being processed, with clear advantages in terms of sample efficiency (Montero-Manso & Hyndman, 2021). Most of the recent successes in applying deep learning to time series forecasting are based on the idea of implementing such global models with a neural network (Benidis et al., 2022). With full generality, in the following, we will consider models belonging to the family in Eq. 2, as Eq. 1 can be seen as a special case where $|{\mathcal{S}}|=1$. In particular, we consider the whole collection at once (i.e., $|{\mathcal{S}}|=N$) and drop the superscript ${\mathcal{S}}$ to simplify the notation.
Template architecture
We use a template architecture analogous to that of Cini et al. (2023b). It consists of the following three processing steps: an encoder, one or more propagation layers, and a decoder. Predictions at each time step and time series are obtained as
${\bm{h}}_{t}^{i,0}=\textsc{Encoder}\left({\bm{x}}_{t-1}^{i},\,{\bm{u}}_{t-1}^{i}\right)$ (3)
${\bm{H}}_{t}^{l+1}=\textsc{PropLayer}^{l}\left({\bm{H}}_{<t}^{l}\right),\quad l=0,1,...,L-1$ (4)
$\mathbf{\hat{x}}_{t:t+H}^{i}=\textsc{Decoder}\left({\bm{h}}_{t}^{i,L},{\bm{u}}_{t:t+H}\right)$ (5)
where the additional index $l$ refers to the layer depth, ${\bm{h}}^{i,l}_{t}\in{\mathbb{R}}^{d_{h}}$ is the representation associated with the $i$-th time series at the $l$-th layer, and ${\bm{H}}^{l}_{t}\in{\mathbb{R}}^{|{\mathcal{S}}|\times d_{h}}$ the stack of such representations. The $\textsc{Encoder}\left(\cdot\right)$ and $\textsc{Decoder}\left(\cdot\right)$ blocks can be implemented in different ways (e.g. a linear layer or a multilayer perceptron (MLP)) and operate w.r.t. single time steps and time series. Conversely, the $\textsc{PropLayer}\left(\cdot\right)$ blocks are the only compontens of the architecture propagating representations across the temporal dimension and across time series in the collection. Each $\textsc{PropLayer}\left(\cdot\right)$ can be implemented by any existing sequence modeling architecture, e.g., RNNs (Hochreiter & Schmidhuber, 1997; Cho et al., 2014a) or temporal convolutional networks (TCNs) (LeCun & Bengio, 1998; Bai et al., 2018), and/or spatiotemporal operators, e.g., spatiotemporal attention (STAtt) models (Grigsby et al., 2021; Deihim et al., 2023). In particular, STGNNs (Seo et al., 2018; Yu et al., 2018), i.e., models that combine sequence modeling architectures with message passing operators, are among the most popular architectures for implementing propagation layers when processing correlated time series (Jin et al., 2023).
Hybrid global-local architectures and local embeddings
As anticipated in the introduction, despite their advantages, global models might struggle to account for the specific dynamics of each time series. Hybrid global-local architectures (Benidis et al., 2022) address this issue by including parameters specific to each target time series. If we indicate these parameters as $\Phi=\{\phi^{0},...,\phi^{N}\}$, the resulting model family would provide forecasts for the time series in the collection as
$\widehat{\bm{X}}^{\mathcal{S}}_{t:t+H}=F\left({\bm{X}}^{\mathcal{S}}_{t-W:t},{\bm{U}}^{\mathcal{S}}_{t-W:t+H};\theta,\Phi^{\mathcal{S}}\right).$ (6)
Although many implementations of such models exist (Benidis et al., 2022; Cini et al., 2023b), we consider models where the local components are implemented as embeddings ${\bm{E}}\in{\mathbb{R}}^{N\times d_{e}}$ of learnable parameters such that $\Phi={\bm{E}}$ and $\phi^{i}={\bm{e}}^{i}$. In particular, each local embedding $\mathbf{e}^{i}$ is associated with the corresponding $i$-th time series and can be learned end-to-end jointly with the shared network weights. Embeddings are incorporated into the template architecture at both the encoder and decoder level by concatenating them to the input as
${\bm{h}}_{t}^{i,0}=\textsc{Encoder}\left({\bm{x}}_{t-1}^{i}\|\,\mathbf{e}^{i},{\bm{u}}_{t-1}^{i}\right),\quad\mathbf{\hat{x}}_{t:t+H}^{i}=\textsc{Decoder}\left({\bm{h}}_{t}^{i,L}\|\,\mathbf{e}^{i},{\bm{u}}_{t:t+H}\right).$ (7)
Fig. 1 provides an overview of the resulting reference architecture. The addition of these parameters comes at a cost in terms of flexibility, as processing time series that were not observed at training time requires fitting new parameters (Januschowski et al., 2020; Cini et al., 2023b).
3 Related works
Learnable embeddings are key components of state-of-the-art time series processing architectures such as STGNNs (Bai et al., 2020; Cini et al., 2023a) and attention-based models (Marisca et al., 2022; Liu et al., 2023; Xiao et al., 2024). In particular, Cini et al. (2023b) systematically addresses the role of such embeddings in hybrid global-local STGNNs. Aside from modeling local dynamics, embeddings have been used extensively to amortize the cost of learning the full adjacency matrix in graph-based models (Wu et al., 2019; Shang et al., 2021; Satorras et al., 2022; De Felice et al., 2024), or as spatial positional encodings (Marisca et al., 2022; Shao et al., 2022; Liu et al., 2023). They are also routinely used in NLP as word embeddings (Mikolov et al., 2013) and, often, as spatial positional encodings (Devlin et al., 2019; Yang et al., 2019).
Methods to regularize learning architectures are obviously central to deep and machine learning in general (Ying, 2019; Tian & Zhang, 2022). Traditional examples include L1 (lasso) (Tibshirani, 1996) and L2 (ridge, weight decay) (Krogh & Hertz, 1991) regularizations. Established approaches in deep learning encompass dropout (Srivastava et al., 2014), layer normalization (Ba et al., 2016) and weight normalization (Salimans & Kingma, 2016). Notably, several regularizations have been tailored to specific architectures (Zaremba et al., 2014; Gal & Ghahramani, 2016; Dieng et al., 2018; Wang & Niepert, 2019; Santos & Papa, 2022), and designed to improve transferability (Wang et al., 2018; Takada & Fujisawa, 2020; Abuduweili et al., 2021). Similarly in spirit to our work, Peng et al. (2015) presents a comparative study of different regularizations for word embeddings in NLP. Regarding regularizations for learnable embeddings for time series, Yin et al. (2022) uses learned embedding clusters to regularize fine-tuning on target data, facilitating transfer in graph-based forecasting. Similarly, Cini et al. (2023b) uses a cluster-based regularization, as well as a variational-based approach. However, these works only address model transferability, without considering the impact of local embedding regularization in a broader context, e.g., transductive settings.
4 Regularization strategies for local embeddings
In this section, we discuss possible shortcomings of hybrid global-local forecasting architectures and present the regularization strategies employed in our experimental analysis.
Figure 2: Validation curves when regularizing whole model or just local parameters in a time series forecasting task (50 runs, ±$\pm 1\sigma$).
Global models are inherently less likely to overspecialize to individual sequences as, having limited capacity, it would compromise performance w.r.t. other sequences within the set. The introduction of local embeddings reduces the regularization effect brought by this inductive bias, resulting in a model that may become more susceptible to overfitting to individual time series, with a potential negative impact on both performance and transferability (Cini et al., 2023b). Model regularization is a common approach to deal with overfitting and to control model and sample complexity. For instance, one might apply well-known techniques, such as weight decay (Krogh & Hertz, 1991) or dropout (Srivastava et al., 2014), across the entire network. However, in many scenarios, the global processing blocks are not particularly overparametrized, as weights are shared. Conversely, in global-local models, one could consider regularizing the local parameters only, i.e., in our settings, constraining how embeddings are learned. Fig. 2 empirically supports this intuition with an example in a simple setting: given a global time-series forecasting model with fixed hyperparameters (green), adding local embeddings (orange) improves performance, but makes it more likely to overfit (as model complexity increases). Regularizing the entire global-local model (purple), however, hinders performance in this scenario. Conversely, applying regularization (dropout in this case) to the embeddings only (red), regularizes the model effectively. Note that regularization of the global parameters can obviously still be beneficial in the case of an over-parametrized global model.
4.1 Regularization methods
In the following, we present the regularization methods that will later be part of our experimental analysis in Sec. 5. They range from adaptations of standard approaches to recent contributions from the literature. We also consider an approach based on parameter resetting.
L1 and L2 regulariation
L2 regularization (Krogh & Hertz, 1991), also known as weight decay, consists in adding to the loss a penalty term proportional to the magnitude of the model’s weights. When applied to the embeddings, the penalty term is $\mathcal{L}_{l2}\left({\bm{E}}\right)=\lambda_{l2}\cdot\|\mathbf{E}\|_{2}^{2}$. Similarly, L1 regularization (Tibshirani, 1996), also known as lasso, consists in applying a penalty term of $\mathcal{L}_{l1}\left({\bm{E}}\right)=\lambda_{l1}\cdot\|\mathbf{E}\|_{1}$. The positive scalar values, $\lambda_{l2}$ and $\lambda_{l1}$, control the regularization strength.
Dropout
Dropout (Srivastava et al., 2014) is another widely recognized regularization. This consists in randomly masking out individual neurons during training, with probability $p$, while scaling the others by $\frac{1}{1-p}$. This technique heuristically forces the network to learn robust representations, often preventing overfitting. In our context, we apply it to the embedding vectors by randomly masking out some parameters for each embedding.
Clustering
The clustering regularization introduced in (Cini et al., 2023b) is based on learning a set of cluster centroids and a cluster assignment matrix, alongside the embeddings. A regularization term is then added to the loss to minimize the distance between embeddings and the assigned centroid. It was originally designed to structure the embedding space and improve model transferability.
Variational regularization
Variational regularization (Cini et al., 2023b), consists in modeling embeddings as samples from a Gaussian posterior learned from data $\mathbf{e}^{i}\sim\mathcal{N}\left(\bm{\mu}^{i},diag\left(\bm{\sigma}^{i}\right)\right)$, where $\bm{\mu}^{i}$ and $\bm{\sigma}^{i}$ are the learnable local parameters. A penalty term based on the KL-divergence between the learned distribution and a standard Gaussian prior, is added to the loss. Similarly to the clustering regularization, it was introduced to enhance model transferability.
Forgetting
We also consider a novel forgetting regularization, based on the forget-and-relearn paradigm (Zhou et al., 2021). This refers to strategies where the parameters of some neural network layers are occasionally reset during training. By doing this, it is possible to reduce memorization of training samples (Baldock et al., 2021), avoiding shortcuts and learning representations that generalize better (Geirhos et al., 2020; Zhou et al., 2021). In practice, we periodically reset the local embeddings ${\bm{E}}$ to a sample from a shared initialization distribution $\mathcal{P}_{e}$ (Narkhede et al., 2022), every $K$ training epochs. Concurrently, we similarly reset the Encoder’s  and Decoder’s weights (Eq. 7) that directly multiply the embeddings (see Appendix C.1). The value of $K$ can be easily selected empirically (see Appendix C.2). Additionally, resetting is halted after a certain amount of epochs to allow for convergence to a final configuration. This can be tuned manually or automatically, by monitoring the validation loss.
4.2 Preventing co-adaptation
As mentioned in Sec. 1, jointly learning local and global parameters can result in their co-adaptation. In particular, embeddings might become simple sequence identifiers and result in models that rely entirely on this identification mechanism. Besides the negative effects on sample efficiency, the resulting architecture would likely lose flexibility in terms of the transferability of the learned representations and processing blocks. In the empirical analysis, we will then focus on whether methods designed to break this co-adaptation can be effective regularizations. Intuitively, the global processing block would be less likely to rely on embeddings as identifiers if these are actively perturbed during training, e.g., by resampling or zeroing.
Among the regularization methods considered in the analysis, L1 and L2 regularizations simply penalize the embeddings’ magnitude. Similarly, clustering encourages embeddings to occupy specific regions within the embedding space and does not actively perturb their values. Conversely, by randomly zeroing out embeddings’ parameters, we might expect dropout to prevent co-adaptation more effectively, as similarly happens when applying it to the weights of subsequent layers (Srivastava et al., 2014). The variational regularization can have a similar impact, as embeddings are constantly resampled during training. Finally, forgetting can prevent co-adaptation in the early stages of training, by periodically resetting the local parameters. Empirical evidence of the effectiveness of such design principles is provided in the following.
5 Experiments
We evaluate the effectiveness of different regularization strategies for local embeddings under three different scenarios: time series forecasting benchmarks (Sec. 5.1), transfer learning (Sec. 5.3), and a sensitivity analysis through embedding perturbations (Sec. 5.4). We consider six real-world datasets of time series collections, spanning four different application domains: METR-LA and PEMS-BAY (Li et al., 2018) as two established benchmarks for traffic forecasting, AQI (Zheng et al., 2015) from the air quality domain, CER-E (CER, 2016) from the energy consumption domain, CLM-D (De Felice et al., 2024) and EngRAD (Marisca et al., 2024) as two multivariate climatic datasets. Details on the datasets, data splits and forecasting settings can be found in Appendix A.
Regarding the investigated models, we consider three different hybrid global-local architectures (Sec. 2.2) distinguished by three different implementations of the propagation layer (Eq. 4). In particular: 1) a RNN with gated recurrent units (GRUs) cells (Cho et al., 2014b) (RNN), 2) a STGNN stacking GRU and anisotropic message-passing layers (Bresson & Laurent, 2017) (STGNN), and 3) a GRU followed by multi-head attention across sequences (Vaswani et al., 2017) (STAtt). Such hybrid architectures are representative of the current state of the art in the considered benchmarks and of the dominant deep learning frameworks for processing sets of related time series (Benidis et al., 2022; Cini et al., 2023a). Implementation and architectural details are provided in Appendix B, while details on the experimental settings can be found in Appendix D.
5.1 Time series forecasting benchmarks
Table 1: Forecasting test error under optimal model size and learning rate (5 runs, ±$\pm 1\sigma$). Methods equal or better than the corresponding standard architectures (RNN, STGNN, STAtt) with embeddings (+Emb.) are in bold. The best-performing method within each dataset and model is in red.
Dataset METR-LA PEMS-BAY CER-E AQI CLM-D EngRAD
Model MAE $\downarrow$ MAE $\downarrow$ MAE $\downarrow$ MAE $\downarrow$ MMRE $\downarrow$ MMRE $\downarrow$
RNN $3.556_{\pm.004}$ $1.774_{\pm.002}$ $0.4453_{\pm.0010}$ $13.279_{\pm.042}$ $19.66_{\pm.01}$ $31.04_{\pm.04}$
+ Emb. $3.148_{\pm.011}$ $1.593_{\pm.004}$ $0.4146_{\pm.0025}$ $13.247_{\pm.044}$ $19.41_{\pm.01}$ $31.01_{\pm.10}$
+ L1 $3.149_{\pm.007}$ $\mathbf{1.590}_{\pm.002}$ $\mathbf{0.4079}_{\pm.0023}$ $\mathbf{13.147}_{\pm.047}$ $19.47_{\pm.01}$ $31.02_{\pm.11}$
+ L2 $\mathbf{3.146}_{\pm.009}$ $\mathbf{1.586}_{\pm.005}$ $\mathbf{0.4058}_{\pm.0011}$ $\mathbf{13.181}_{\pm.026}$ $19.42_{\pm.01}$ $\mathbf{30.97}_{\pm.06}$
+ Clust. $\mathbf{3.138}_{\pm.015}$ $\mathbf{1.588}_{\pm.007}$ $\mathbf{0.4115}_{\pm.0025}$ $\mathbf{13.231}_{\pm.044}$ $\mathbf{19.41}_{\pm.01}$ $31.05_{\pm.07}$
+ Drop. $\mathbf{3.147}_{\pm.012}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{1.580}}_{\pm.007}$ $\mathbf{0.4104}_{\pm.0005}$ $\mathbf{13.114}_{\pm.038}$ $19.43_{\pm.01}$ $\mathbf{30.99}_{\pm.09}$
+ Vari. ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{3.132}}_{\pm.006}$ $\mathbf{1.589}_{\pm.006}$ $\mathbf{0.4050}_{\pm.0006}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{13.113}}_{\pm.029}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{19.39}}_{\pm.00}$ $\mathbf{30.98}_{\pm.05}$
+ Forg. $3.149_{\pm.007}$ $\mathbf{1.590}_{\pm.007}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{0.4049}}_{\pm.0007}$ $\mathbf{13.185}_{\pm.022}$ $19.42_{\pm.02}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{30.92}}_{\pm.02}$
STGNN $3.239_{\pm.017}$ $1.660_{\pm.003}$ $0.4275_{\pm.0006}$ $11.814_{\pm.051}$ $19.19_{\pm.03}$ $28.04_{\pm.08}$
+ Emb. $3.027_{\pm.009}$ $1.593_{\pm.004}$ $0.4144_{\pm.0032}$ $11.881_{\pm.053}$ $18.89_{\pm.04}$ $27.52_{\pm.09}$
+ L1 $3.040_{\pm.016}$ $\mathbf{1.587}_{\pm.005}$ $\mathbf{0.4039}_{\pm.0009}$ $\mathbf{11.789}_{\pm.043}$ $18.92_{\pm.03}$ $\mathbf{27.45}_{\pm.12}$
+ L2 $\mathbf{3.023}_{\pm.009}$ $\mathbf{1.582}_{\pm.003}$ $\mathbf{0.4016}_{\pm.0014}$ $\mathbf{11.795}_{\pm.025}$ $\mathbf{18.87}_{\pm.04}$ $\mathbf{27.44}_{\pm.14}$
+ Clust. $\mathbf{3.025}_{\pm.012}$ $\mathbf{1.580}_{\pm.005}$ $\mathbf{0.4075}_{\pm.0020}$ $\mathbf{11.876}_{\pm.053}$ $\mathbf{18.85}_{\pm.04}$ $27.60_{\pm.14}$
+ Drop. $3.036_{\pm.011}$ $\mathbf{1.575}_{\pm.006}$ $\mathbf{0.4042}_{\pm.0008}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{11.712}}_{\pm.016}$ $18.93_{\pm.04}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{27.41}}_{\pm.06}$
+ Vari. ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{3.013}}_{\pm.005}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{1.566}}_{\pm.003}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{0.3989}}_{\pm.0012}$ $\mathbf{11.768}_{\pm.026}$ $\mathbf{18.85}_{\pm.02}$ $27.53_{\pm.10}$
+ Forg. $3.050_{\pm.017}$ $\mathbf{1.578}_{\pm.006}$ $\mathbf{0.4026}_{\pm.0006}$ $\mathbf{11.793}_{\pm.040}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{18.83}}_{\pm.02}$ $\mathbf{27.47}_{\pm.12}$
STAtt $3.538_{\pm.004}$ $1.776_{\pm.002}$ $0.4479_{\pm.0014}$ $13.341_{\pm.330}$ $19.74_{\pm.02}$ $29.25_{\pm.09}$
+ Emb. $3.074_{\pm.020}$ $1.616_{\pm.012}$ $0.4143_{\pm.0022}$ $12.973_{\pm.402}$ $19.10_{\pm.03}$ $28.32_{\pm.08}$
+ L1 $\mathbf{3.061}_{\pm.021}$ $\mathbf{1.590}_{\pm.005}$ $\mathbf{0.4072}_{\pm.0009}$ $\mathbf{12.799}_{\pm.697}$ $19.27_{\pm.04}$ $\mathbf{28.15}_{\pm.33}$
+ L2 $\mathbf{3.058}_{\pm.017}$ $\mathbf{1.593}_{\pm.004}$ $\mathbf{0.4061}_{\pm.0004}$ $13.542_{\pm.606}$ $19.17_{\pm.06}$ $\mathbf{28.12}_{\pm.19}$
+ Clust. $\mathbf{3.061}_{\pm.021}$ $\mathbf{1.590}_{\pm.008}$ $\mathbf{0.4084}_{\pm.0017}$ $\mathbf{12.535}_{\pm.240}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{19.06}}_{\pm.01}$ $\mathbf{27.96}_{\pm.17}$
+ Drop. $\mathbf{3.058}_{\pm.005}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{1.565}}_{\pm.004}$ $\mathbf{0.4068}_{\pm.0012}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{12.443}}_{\pm.482}$ $19.30_{\pm.08}$ $\mathbf{28.12}_{\pm.23}$
+ Vari. ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{3.041}}_{\pm.009}$ $\mathbf{1.571}_{\pm.005}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{0.4030}}_{\pm.0014}$ $\mathbf{12.616}_{\pm 1.145}$ $\mathbf{19.10}_{\pm.06}$ $\mathbf{28.01}_{\pm.13}$
+ Forg. $\mathbf{3.058}_{\pm.011}$ $\mathbf{1.579}_{\pm.003}$ $\mathbf{0.4062}_{\pm.0012}$ $\mathbf{12.455}_{\pm.416}$ $19.16_{\pm.02}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{27.94}}_{\pm.14}$
In our first experiment, we consider the problem of time series forecasting in a transductive setting, i.e., the set of time series to be forecast is the same set observed at training time.
Table 2: Summary statistics for Tab. 1. Columns indicate the average improvement (%) over the reference un-regularized model and mean relative rank. Best is in red, second in bold.
Reg. %impr. Rank
L1 $0.58$ $4.6$
L2 $0.45$ $3.4$
Clust. $0.67$ $4.0$
Drop. $0.94$ $3.2$
Vari. ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{1.21}}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{2.0}}$
Forg. $\mathbf{0.95}$ $\mathbf{3.0}$
The forecasting horizon is dataset-dependent and reported in Tab. 4 of Appendix A. We use models without local parameters and with un-regularized local embeddings as reference baselines. Then, we evaluate performance with regularized local embeddings, adopting the different strategies detailed in Sec. 4.1. For model selection, a hyperparameter search on model size and learning rate has been carried out independently for each model variant and dataset. All regularization hyperparameters have been set as detailed in Appendix D. Tab. 1 shows the obtained results, while Tab. 2 provides a concise summary of the performance of each regularization across experimental settings. Regularizing the learning of local embeddings provides performance improvements over non-regularized models in most datasets (F1). Considering the negligible computational overhead and ease of implementation, these results support the adoption of such techniques as standard practice. While there is no clear winner among the different regularization methods, the variational regularization appears to be the most effective, on average; followed by forgetting and dropout. This supports adopting methods that prevent co-adaptation by design, i.e., that prevent the global components from relying on specific values of the local parameters (F2) (F3). Additional results on the impact of regularizing embeddings when Eq. 4 is implemented by a simple MLP are reported in Appendix E.
5.2 Sensitivity analysis and learning curves
Figure 3: Validation curves for different training scenarios (5 runs, ±$\pm 1\sigma$). Plot names follow the convention model (embedding size) [dataset].
To provide additional insight, we investigate how different regularizations affect the learning curve across different embedding sizes. In doing so, we fix the shared model hidden size to $h=64$ and learning rate to $\eta=0.00075$. Specifically, Fig. 3 shows the validation mean absolute error (MAE) across training epochs for some example scenarios; different colors correspond to different regularization strategies. For completeness, additional plots in complementary settings are reported in Appendix F. Comparing the different regularizations, we can see that dropout (red) and forgetting (blue) significantly affect the learning curves, even with large embeddings (Fig. 3, right). However, dropout, as one could expect, might be problematic when applied to embeddings of limited size (Fig. 3, left). The forgetting regularization is robust across the different configurations. The variational regularization appears less disruptive of the learning dynamics. The remaining regularizations (i.e., L1, L2, and clustering) have little impact on the learning curve. Overall, regularization strategies that perturb the local embeddings seem more likely to positively affect the learning curves, which might be a reason for the superior performance shown in Tab. 2.
5.3 Transfer learning
Table 3: Forecasting test error (MAE) in transfer learning (5 runs, ±$\pm 1\sigma$, STGNN). Within each target dataset, different rows pertain to different fine-tuning data lengths. Bold denotes methods equal or better than the baseline (+ Emb.), best method is in red.
Tr. size + Emb. + L1 + L2 + Clust. + Dropout + Varia. + Forget.
PEMS03 Zero-shot $26.19_{\pm.98}$ $\mathbf{26.06}_{\pm.61}$ $\mathbf{25.26}_{\pm.68}$ $\mathbf{25.18}_{\pm.40}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{21.45}}_{\pm.26}$ $\mathbf{22.71}_{\pm.13}$ $\mathbf{25.08}_{\pm.46}$
1 day $18.92_{\pm.08}$ $\mathbf{18.89}_{\pm.10}$ $\mathbf{18.87}_{\pm.08}$ $\mathbf{18.77}_{\pm.05}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{17.99}}_{\pm.02}$ $\mathbf{18.17}_{\pm.07}$ $\mathbf{18.55}_{\pm.09}$
3 days $18.41_{\pm.04}$ $\mathbf{18.41}_{\pm.07}$ $\mathbf{18.35}_{\pm.05}$ $18.53_{\pm.06}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{17.86}}_{\pm.05}$ $\mathbf{17.91}_{\pm.03}$ $\mathbf{18.19}_{\pm.03}$
1 week $17.53_{\pm.02}$ $\mathbf{17.52}_{\pm.05}$ $\mathbf{17.47}_{\pm.07}$ $17.59_{\pm.06}$ $\mathbf{17.31}_{\pm.04}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{17.26}}_{\pm.02}$ $\mathbf{17.34}_{\pm.03}$
2 weeks $17.34_{\pm.03}$ $\mathbf{17.34}_{\pm.03}$ $\mathbf{17.28}_{\pm.04}$ $17.37_{\pm.04}$ $\mathbf{17.20}_{\pm.02}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{17.16}}_{\pm.02}$ $\mathbf{17.20}_{\pm.05}$
PEMS04 Zero-shot $29.23_{\pm.43}$ $\mathbf{29.09}_{\pm.26}$ $29.26_{\pm.47}$ $30.03_{\pm.26}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{26.13}}_{\pm.31}$ $\mathbf{27.56}_{\pm.31}$ $\mathbf{28.44}_{\pm.34}$
1 day $23.93_{\pm.16}$ $24.00_{\pm.10}$ $23.96_{\pm.16}$ $\mathbf{23.71}_{\pm.05}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{23.04}}_{\pm.04}$ $\mathbf{23.33}_{\pm.12}$ $\mathbf{23.54}_{\pm.13}$
3 days $23.18_{\pm.12}$ $23.22_{\pm.14}$ $23.27_{\pm.10}$ $\mathbf{22.97}_{\pm.06}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{22.67}}_{\pm.03}$ $\mathbf{22.87}_{\pm.06}$ $\mathbf{22.86}_{\pm.08}$
1 week $22.47_{\pm.07}$ $\mathbf{22.46}_{\pm.05}$ $\mathbf{22.45}_{\pm.07}$ $22.49_{\pm.07}$ $\mathbf{22.29}_{\pm.03}$ $\mathbf{22.38}_{\pm.05}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{22.26}}_{\pm.03}$
2 weeks $21.92_{\pm.04}$ $21.93_{\pm.04}$ $21.93_{\pm.05}$ $22.00_{\pm.01}$ $21.96_{\pm.03}$ $22.04_{\pm.04}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{21.79}}_{\pm.03}$
PEMS07 Zero-shot $57.40_{\pm 6.70}$ $\mathbf{55.30}_{\pm 3.97}$ $\mathbf{53.97}_{\pm 3.37}$ $\mathbf{56.92}_{\pm 1.62}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{34.19}}_{\pm.78}$ $\mathbf{42.99}_{\pm 2.83}$ $\mathbf{42.67}_{\pm 2.64}$
1 day $29.61_{\pm.17}$ $29.63_{\pm.24}$ $\mathbf{29.37}_{\pm.27}$ $\mathbf{29.00}_{\pm.32}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{27.25}}_{\pm.12}$ $\mathbf{28.30}_{\pm.41}$ $\mathbf{28.38}_{\pm.19}$
3 days $27.65_{\pm.14}$ $27.74_{\pm.14}$ $\mathbf{27.56}_{\pm.14}$ $\mathbf{27.56}_{\pm.14}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{26.37}}_{\pm.10}$ $\mathbf{26.92}_{\pm.21}$ $\mathbf{26.92}_{\pm.08}$
1 week $26.60_{\pm.04}$ $26.66_{\pm.09}$ $\mathbf{26.60}_{\pm.07}$ $26.90_{\pm.14}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{25.87}}_{\pm.13}$ $\mathbf{26.20}_{\pm.17}$ $\mathbf{26.03}_{\pm.06}$
2 weeks $25.84_{\pm.04}$ $25.90_{\pm.07}$ $\mathbf{25.76}_{\pm.04}$ $26.31_{\pm.24}$ $\mathbf{25.53}_{\pm.15}$ $\mathbf{25.66}_{\pm.17}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{25.41}}_{\pm.04}$
PEMS08 Zero-shot $25.41_{\pm.62}$ $\mathbf{25.26}_{\pm 1.09}$ $25.90_{\pm.47}$ $25.60_{\pm.44}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{21.04}}_{\pm.17}$ $\mathbf{22.60}_{\pm.37}$ $\mathbf{23.90}_{\pm.43}$
1 day $18.91_{\pm.09}$ $\mathbf{18.88}_{\pm.05}$ $\mathbf{18.85}_{\pm.08}$ $\mathbf{18.64}_{\pm.08}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{17.96}}_{\pm.06}$ $\mathbf{18.23}_{\pm.12}$ $\mathbf{18.40}_{\pm.13}$
3 days $18.11_{\pm.10}$ $18.16_{\pm.13}$ $18.15_{\pm.08}$ $\mathbf{18.03}_{\pm.04}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{17.55}}_{\pm.06}$ $\mathbf{17.77}_{\pm.07}$ $\mathbf{17.83}_{\pm.14}$
1 week $17.33_{\pm.08}$ $\mathbf{17.28}_{\pm.06}$ $\mathbf{17.33}_{\pm.03}$ $17.40_{\pm.04}$ $\mathbf{17.22}_{\pm.06}$ $\mathbf{17.31}_{\pm.07}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{17.11}}_{\pm.06}$
2 weeks $17.13_{\pm.06}$ $17.18_{\pm.14}$ $17.15_{\pm.07}$ $17.25_{\pm.03}$ $17.15_{\pm.08}$ $17.20_{\pm.09}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{16.99}}_{\pm.05}$
Our second main experiment consists of time series forecasting under a transfer learning setting, adapted from previous works (Cini et al., 2023b). This benchmark aims to verify the impact of regularizing the local embeddings on the transferability of the global (shared) processing blocks. We consider the $4$ PEMS benchmark datasets from Guo et al. (2021) (i.e., PEMS03, PEMS04, PEMS07, PEMS08) and a reference STGNN model with local embeddings. For each subset of three datasets, we first train the entire model, then we reset the embeddings to their initial values and fine-tune them exclusively on the held-out dataset. During fine-tuning, shared parameters are kept frozen. The different regularizations are applied only during the initial training procedure and disabled during fine-tuning. Tab. 3 reports the results for the un-regularized and regularized models as the amount of data for fine-tuning varies (zero-shot refers to no fine-tuning at all). It is evident how dropout, variational regularization, and forgetting are the most reliable and effective strategies (F2). Notably, dropout excels in the zero-shot setting, while forgetting is consistent across fine-tuning lengths and excels when more data are available (F3). This suggests dropout can be particularly useful when data from the transfer domain is scarce or absent.
5.4 Robustness to local parameter perturbation
Figure 4: Test performance degradation on embeddings perturbation (5 runs, ±$\pm 1\sigma$, STGNN, METR-LA). Left: Adding zero-mean Gaussian noise. Right: (Left) random shuffling, (Middle) replaced with their mean, and (Right) replaced by a draw from their sample normal.
Finally, we investigate how a model’s forecasting accuracy is affected by the perturbation of the learned local embeddings. This provides insights into the robustness of the shared learned parameters and can serve as a proxy for the effectiveness of different regularizations in preventing co-adaptation. To avoid penalizing regularizations that are sensitive to the weights’ magnitude  (i.e., L2, L1), we consider perturbations that do not impact the scale of the learned representations. In particular, we experiment with four such strategies: adding zero-mean Gaussian noise to the embeddings (Noise $\sigma$), randomly shuffling embeddings across sequences (Rearranged), replacing each embedding with their sample mean (Mean), resampling each embedding from their sample normal distribution (Sampled) (see Appendix D.1 for a formal definition). Fig. 4 shows the results of the analysis after training an STGNN on the METR-LA dataset (complementary settings are illustrated in Appendix F). Regularization methods that actively perturb the embeddings while learning (i.e., dropout, forgetting, variational) consistently result in more robust models under all the considered perturbations (F2) (F3). Conversely, other regularizations have marginal, or even negative, impact. We observe some consistency between the results observed in Fig. 4 and performance in transfer learning, shown in Tab. 3.
6 Conclusions
This paper highlights the importance of regularizing the learning of local embeddings in modern deep-learning architectures for time series forecasting. Our empirical study, across diverse datasets and scenarios, provides clear evidence that this practice is beneficial for a variety of reference architectures, representative of the state of the art. Notably, even simple techniques, when applied to local embeddings, can yield consistent performance gains. Furthermore, we observed that methods that actively perturb the embeddings at training time, such as dropout, variational regularization, and forgetting, consistently rank among the top performers, both in transductive and transfer learning settings. Overall, our findings offer a strong empirical argument for the adoption of embedding regularization as standard practice when designing hybrid global-local architectures for time series. We believe this is an important building block towards developing more robust and transferable models for spatiotemporal data and designing foundation models for related time series processing.
Limitations and future works
While it is clear that these regularizations provide consistent advantages, it is difficult to strongly identify the best-performing method in the different scenarios. Hence, future works might focus on finding strategies that combine the best qualities of the different methods. Furthermore, future studies could try to provide methods to quantitatively and analytically characterize the co-adaptation of global and local parameters, for which we can only identify indirect signs. In particular, they might design experiments to detect the degeneration of local components into mere identifiers. Moreover, future research could explore how the efficacy of different regularization methods varies w.r.t. the number of input time series and different downstream tasks.
Appendix
Appendix A Datasets
DATASETS # Time series Time steps Channels Connectivity Edges Sampling Rate Time window Horizon
METR-LA 207 34,272 1 Directed 1515 5 minutes 12 12
PEMS-BAY 325 52,128 1 Directed 2369 5 minutes 12 12
CER-E 485 25,728 1 Directed 4365 30 minutes 48 6
AQI 437 8,760 1 Undirected 2699 1 hour 24 3
CLM-D 235 10,958 10 NA 2699 1 day 14 3
EngRAD 487 26,304 5 NA 2699 1 hour 24 6
PEMS03 358 26,208 1 Directed 546 5 minutes 12 12
PEMS04 307 16,992 1 Directed 340 5 minutes 12 12
PEMS07 883 28,224 1 Directed 866 5 minutes 12 12
PEMS08 170 17,856 1 Directed 277 5 minutes 12 12
Table 4: Dataset details.
METR-LA
traffic data from Li et al. (2018). Traffic readings are from different loop detectors on highways in the Los Angeles County. Licensed under Attribution 4.0 International (CC BY 4.0).
PEMS-BAY
traffic data from Li et al. (2018). Traffic readings are from different loop detectors on highways in the Los Angeles County. Licensed under Attribution 4.0 International (CC BY 4.0).
CER-E
Electric load data from (CER, 2016). Data encompass energy consumption readings from smart meters in small and medium enterprises, collected in the context of the Commission for Energy Regulation (CER) Smart Metering Project. Data access can be requested through https://www.ucd.ie/issda/data/commissionforenergyregulationcer/
AQI
Air quality data from Zheng et al. (2015). The dataset collects measurements of the PM2.5 pollutant from air quality stations in 43 Chinese cities and is available at https://www.microsoft.com/en-us/research/publication/forecasting-fine-grained-air-quality-based-on-big-data/
CLM-D
[…]
EngRAD
[…]
PEMS03, PEMS04, PEMS07, and PEMS08
datasets from  Guo et al. (2021) collect traffic detector data from 4 districts in California provided by Caltrans Performance Measurement System (PeMS). Data are aggregated into 5-minutes intervals.
All datasets were split 70%/10%/20% into train, validation and test along the temporal axis. Datasets with a high number of missing values, i.e., METR-LA, PEMS-BAY and AQI, have a binary mask concatenated to the input, indicating if the corresponding value has been artificially imputed. Note that, unless differently specified, used datasets are public domain. Where possible, we pointed to the original source of the data or to a meaningful reference.
Appendix B Models
In this section we describe the models used in our study. Note that all models used a fixed hidden size $d_{h}$ for all layers. For all the architectures, the Encoder 3 is parametrized by a linear layer, while the Decoder is a 1-layer MLP followed by $H$ parallel linear layers, each decoding a different step in the forecasting horizon. The RNN model is implemented by means of a 1-layer GRU (Cho et al., 2014b) shared among all sequences. On top of the same GRU architecture, for the STGNN model, we employ 2 layers of message passing defined as
$\mathbf{m}^{j\rightarrow i}$ $=\mathbf{W}_{2}\xi\left(\mathbf{W}_{1}\left[{\bm{h}}^{i}\|{\bm{h}}^{j}\|a_{ji}\right]\right),\qquad\alpha^{j\rightarrow i}=\sigma\left(\mathbf{W}_{0}\mathbf{m}^{j\rightarrow i}\right),$ (8)
$\mathbf{\tilde{h}}^{i}$ $=\xi\left(\mathbf{W}_{3}{\bm{h}}^{i}+\sum_{j\in\mathcal{N}\left(i\right)}\{\alpha^{j\rightarrow i}\mathbf{m}^{j\rightarrow i}\}\right),$ (9)
[…]
$\mathbf{\tilde{h}}^{i}=\xi\left(\mathbf{W}_{1}{\bm{h}}^{i}+\frac{1}{\|\mathcal{N}\left(i\right)\|}\sum_{j\in\mathcal{N}\left(i\right)}\{\mathbf{W}_{2}{\bm{h}}^{j}\}\right),$ (10)
[…]
Appendix C Additional details on forgetting regularization
For the sake of completeness, in the following, we list some additional details regarding the local forgetting regularization.
C.1 Embedding-related Encoder/Decoder parameters reset
When resetting the local parameters, we also reset the encoder’s (Eq. 3) and decoder’s (Eq. 5) parameters (i.e., coefficients of a linear layer) that directly interact with the embeddings’ features. As an illustrative example, consider an Encoder parametrized by a MLP, with input linear layer
${\bm{h}}_{t}^{0}=\left[{\bm{X}}_{t-1}\|{\bm{U}}_{t-1}\|\mathbf{E}\right]{\bm{W}}^{T}+{\bm{b}},$ (11)
In this case, the encoder’s parameters to be reset would correspond to the last $d_{e}$ columns of the weight matrix ${\bm{W}}$. An equivalent behavior is implemented for the decoder’s input layer.
C.2 Forgetting period sensitivity
(a) METR-LA
(b) PEMS-BAY
(c) AQI
Figure 5: Sensitivity to forgetting period $k$ across different learning rates and datasets (5 runs, STGNN). A period of 00 indicates no regularization.
The introduced forgetting regularization has two hyperparameters: the reset period $k$ and the halting epoch. While the latter can be determined automatically, by monitoring the validation error before each reset, the former should be set empirically. To provide insight on the degree to which the selection of $k$ can impact regularization performance, we perform a sensitivity study on $3$ datasets: METR-LA, PEMS-BAY and AQI. The results are shown in Fig. 5, where we report the mean absolute error (MAE)with varying learning rates and values of $k$. We can see that performance is relatively stable w.r.t. changes in $k$, suggesting that hyperparameter search can be limited to few values (e.g., $3$ for a short period and $30$ for a long one). Another possibility, to avoid searches, is to select the value by hand, after inspecting the learning curves of the unregularized model, in order to select a reset period that allows performance recovery.
C.3 Forgetting warm-up
To avoid instabilities at the beginning of training, particularly when using short reset periods, the forgetting routine can be initiated after a warm-up period, which can be set empirically to a few epochs, as commonly done with learning rate warm-up procedures or with other regularizations (Cini et al., 2023b).
Appendix D Detailed experimental setting
Shared settings
All the models were trained with the Adam optimizer (Kingma & Ba, 2015), with a batch size of 64 and up to 300 batches per epoch. We used the Python (Van Rossum & Drake, 2009) programming language, leveraging Torch Spatiotemporal (Cini & Marisca, 2022), Pytorch (Paszke et al., 2019) and Pytorch Lightning (Falcon & The PyTorch Lightning team, 2019) to implement all the experiments. Experiments were scheduled and logged by leveraging Hydra (Yadan, 2019) and Weights and Biases (Biewald, 2020). The mean absolute error (MAE)was used as loss function in all experiments. For the different regularization’s hyperparameters, we heuristically found sensible values (i.e., with stable performance across datasets, models and other hyperparameters) or used those of previous works, in particular: dropout with probability of dropping a connection $p=0.5$, weight of L2 $\lambda_{l2}=0.0001$, weight of L1 $\lambda_{l1}=0.00001$, resetting period of forgetting $k=20$ epochs with $30$ epochs of warm-up, weight of variational regularization $\lambda_{var}=0.00005$ and weight of clustering regularization $\lambda_{clst}=0.0005$. Notice that, in the transfer experiment 3, for variational and clustering regularizations, we used the values adopted in the original paper (Cini et al., 2023b), $\lambda_{var}=0.05$ and $\lambda_{clst}=0.5$ respectively. Any other hyperparameter of variational and clustering regularizations was taken from the original paper as well. Unless specified, we employed and embedding size $d_{e}=32$ for all models. Moreover, all datasets used temporal encodings as additional covariates, in particular, a one-hot encoding of the weekday and sinus and cosinus encodings with daily period.
Experiment specific hyperparameters
For the experiments in Tab. 1, optimal hyperparameters for each model, regularization and dataset were found via a grid-search over learning rates values in the range $lr\in[0.00025,0.003]$ and model hidden size value $d_{h}\in[32,256]$, with constraints caused by GPU memory capacity in some settings. Forgetting was halted after $150$ epochs. Each model was trained for 200 epochs and best validation parameters were used for testing. As for the transfer learning experiments in Tab. 3, we followed the settings in Cini et al. (2023b), furthermore, we used a learning rate $lr=0.005$ during training and $lr=0.001$ during fine-tuning. In such scenario, the training lasted up to 150 epochs with 50 epochs of early stopping patience, while fine-tuning lasted up to 1000 epochs with 100 epochs of patience. No temporal encodings were used in the transfer setting. For all the other experiments we used an hidden size $d_{h}=64$ and learning rate $lr=0.00075$, with 300 maximum epochs and early stopping set to 100, while forgetting was halted after 100 epochs. Note that forgetting ignores early stopping while active, to avoid triggering it as a byproduct of the embedding resetting itself.
Metrics
For univariate datasets, we consider the mean absolute error (MAE), as this is a standard choice across the related literature. For multivariate datasets, i.e., CLM-D and EngRAD, following De Felice et al. (2024), we compute the mean relative error (MRE)independently for each channel, then, take the channel-wise average to obtain a unique final score, which we termed multivariate mean relative error (MMRE)in Tab. 1.
Computing resources
Experiments were run on A100 and A5000 NVIDIA GPUs. The vast majority of the experiments conducted in our work can be easily run on moderate GPU hardware, with at least 8GBs of VRAM. However, not that some experiments, may require more. For instance, the transfer experiments require at least 20GBs of VRAM on some of the benchmark dataset. Moreover, high hidden size configurations of STGNN and STAtt models, required up to 40GBs of VRAM. In general, a single run (i.e., training one model on one dataset), given hardware that fulfills the memory requirements, takes from 30 minutes to 3 hours, depending on the specific configuration.
D.1 Embeddings perturbation
Here we provide a more formal definition of the perturbations adopted in Sec. 5.4. In particular, considering $N$ embedding vectors:
  • Noise $\sigma$ consists in adding zero-mean gaussian noise, from an isotropic multivariate normal distribution, to the embeddings. Formally, this refers to substituting each embedding $\mathbf{e}^{i}$ with $\mathbf{e}^{i}+\epsilon$, where $\epsilon\sim\mathcal{N}\left(\mathbf{0},diag\left(\sigma^{2}\right)\right)$.
  • Rearrange refers to randomly reassigning the embedding vectors to different time-series in the collection. Formally, this means substituting each embedding $\mathbf{e}^{i}$ with an embedding $\mathbf{e}^{j}$, where $j\sim Multinomial\left(\{0,...,N-1\}\right)$ is a sample from a multinomial distribution over the embedding indices, sampled without repetition.
  • Mean refers to setting each embedding to the mean values across embeddings themselves. Formally, replacing each embedding $\mathbf{e}^{i}$ with $\frac{1}{N}\sum_{j=0}^{N}\mathbf{e}^{j}$.
  • Sampled consists in estimating the sample gaussian distribution of the embeddings, and then replacing each embedding with a draw from such distribution. Formally, this entails estimating the gaussian’s parameters as $\mu_{e}=\frac{1}{N}\sum_{i=0}^{N-1}\mathbf{e}^{i}$ and $\sigma^{2}_{e}=\frac{1}{N-1}\sum_{i=0}^{N-1}\left(\mathbf{e}^{i}-\mu\right)$. Then we sample values for each embedding as $\mathbf{e}^{i}\sim\mathcal{N}\left(\mu_{e},\sigma^{2}_{e}\right)$.
Note that, to avoid penalizing regularizations that are sensible to the weights’ magnitude  (i.e., L1, L2), we consider perturbations that do not change the scale of the learned representations. As this does not hold true for the Noise $\sigma$ perturbation, since it can potentially affect the embeddings magnitude, depending on the standard deviation, we selected most values to be within reason w.r.t. the embeddings magnitude itself. Nonetheless, we also employed a few more extreme values, in order to observe their effects.
Appendix E Effect on simple global models
The purpose of model regularization is, usually, to limit model capacity by means of additional constraints and, in turn, obtain models that generalize better. This implies that excessive regularization might result in a degradation of performance. In principle, regularization of local parameters should not have a negative impact on underparametrized global models, as their parameters are not constrained directly. To evaluate this, we employ a simple 2-layer MLP as our global model (Eq. 4), and train it with the same settings as in Sec. 5.1.
Table 5: Forecasting error in trasductive setting for MLP model (hidden size 64-128) (5 runs, ±$\pm 1\sigma$). Equal/better than unregularized (+Emb) in bold. Best in red.
Dataset METR-LA PEMS-BAY CER-E AQI CLM-D EngRAD
Model MAE $\downarrow$ MAE $\downarrow$ MAE $\downarrow$ MAE $\downarrow$ MMRE $\downarrow$ MMRE $\downarrow$
MLP (64) $3.580_{\pm 0.005}$ $1.808_{\pm 0.002}$ $0.4658_{\pm 0.0008}$ $\mathbf{13.398}_{\pm 0.044}$ $19.86_{\pm 0.01}$ $\mathbf{31.38}_{\pm 0.13}$
+ Emb. $3.153_{\pm 0.012}$ $1.619_{\pm 0.005}$ $0.4233_{\pm 0.0007}$ $13.521_{\pm 0.082}$ $19.55_{\pm 0.01}$ $31.40_{\pm 0.14}$
+ L1 $\mathbf{3.150}_{\pm 0.005}$ $\mathbf{1.614}_{\pm 0.006}$ $0.4236_{\pm 0.0008}$ $\mathbf{13.392}_{\pm 0.067}$ $19.56_{\pm 0.01}$ $\mathbf{31.32}_{\pm 0.04}$
+ L2 $\mathbf{3.147}_{\pm 0.007}$ $\mathbf{1.613}_{\pm 0.004}$ $0.4242_{\pm 0.0010}$ $\mathbf{13.389}_{\pm 0.042}$ $19.57_{\pm 0.02}$ $\mathbf{31.21}_{\pm 0.09}$
+ Clust. $\mathbf{3.152}_{\pm 0.014}$ $1.625_{\pm 0.009}$ $0.4236_{\pm 0.0012}$ $\mathbf{13.478}_{\pm 0.070}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{19.54}}_{\pm 0.01}$ $\mathbf{31.25}_{\pm 0.14}$
+ Drop. $3.193_{\pm 0.003}$ $1.630_{\pm 0.006}$ $0.4317_{\pm 0.0009}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{13.328}}_{\pm 0.036}$ $19.62_{\pm 0.01}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{31.20}}_{\pm 0.07}$
+ Vari. ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{3.139}}_{\pm 0.011}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{1.610}}_{\pm 0.001}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{0.4216}}_{\pm 0.0011}$ $\mathbf{13.337}_{\pm 0.020}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{19.54}}_{\pm 0.01}$ $\mathbf{31.32}_{\pm 0.03}$
+ Forg. $3.157_{\pm 0.013}$ $\mathbf{1.617}_{\pm 0.007}$ $0.4271_{\pm 0.0012}$ $\mathbf{13.472}_{\pm 0.099}$ $19.56_{\pm 0.02}$ $\mathbf{31.34}_{\pm 0.09}$
MLP (128) $3.571_{\pm 0.002}$ $1.798_{\pm 0.003}$ $0.4557_{\pm 0.0003}$ $\mathbf{13.390}_{\pm 0.044}$ $19.78_{\pm 0.01}$ $31.44_{\pm 0.04}$
+ Emb. $3.157_{\pm 0.014}$ $1.615_{\pm 0.006}$ $0.4141_{\pm 0.0013}$ $13.534_{\pm 0.052}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{19.47}}_{\pm 0.01}$ $31.38_{\pm 0.10}$
+ L1 $\mathbf{3.152}_{\pm 0.004}$ $\mathbf{1.604}_{\pm 0.006}$ $\mathbf{0.4128}_{\pm 0.0007}$ $\mathbf{13.417}_{\pm 0.078}$ $19.49_{\pm 0.02}$ $\mathbf{31.38}_{\pm 0.07}$
+ L2 $\mathbf{3.150}_{\pm 0.015}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{1.600}}_{\pm 0.006}$ $\mathbf{0.4117}_{\pm 0.0009}$ $\mathbf{13.450}_{\pm 0.059}$ $19.48_{\pm 0.01}$ $31.65_{\pm 0.17}$
+ Clust. $\mathbf{3.143}_{\pm 0.017}$ $\mathbf{1.603}_{\pm 0.003}$ $\mathbf{0.4134}_{\pm 0.0009}$ $13.612_{\pm 0.119}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{19.47}}_{\pm 0.01}$ $31.42_{\pm 0.06}$
+ Drop. $3.195_{\pm 0.008}$ $1.624_{\pm 0.003}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{0.4101}}_{\pm 0.0003}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{13.320}}_{\pm 0.044}$ $19.53_{\pm 0.02}$ $\mathbf{31.35}_{\pm 0.13}$
+ Vari. ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{3.139}}_{\pm 0.017}$ $\mathbf{1.603}_{\pm 0.005}$ $\mathbf{0.4109}_{\pm 0.0012}$ $\mathbf{13.346}_{\pm 0.038}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{19.47}}_{\pm 0.01}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{31.33}}_{\pm 0.07}$
+ Forg. $\mathbf{3.150}_{\pm 0.014}$ $1.616_{\pm 0.010}$ $0.4184_{\pm 0.0008}$ $\mathbf{13.427}_{\pm 0.058}$ $19.53_{\pm 0.02}$ $\mathbf{31.36}_{\pm 0.09}$
Tab. 5 reports the obtained results in case of tuned learning rate and hidden size of $64$ and $128$ units. Compared to models in Tab. 1, the performance of the MLP is worse or comparable at best, however, the regularizations relative effectiveness is similar, w.r.t. each dataset, at hidden size $128$. In the extreme case of the MLP with hidden size $64$, we can observe additional scenarios in which most regularizations are ineffective (i.e., CER-E). Nonetheless, we do not observe catastrophic effects on performance. Notably, with some datasets, i.e., AQI and EngRAD, adding local parameters to the global MLP model hurts performance, while regularization allows an effective exploitation of the embeddings.
Appendix F Additional experiments
In this section, we provide additional experimental results, for completeness w.r.t. what has been shown in the main paper.
Figure 6: Validation curves for different model families (5 runs, ±$\pm 1\sigma$). Curves show the application of dropout regularization at the embedding level or over the entire architecture. Plot names follow the convention model (embedding size) [dataset].
In particular, Fig. 6 complements what has been shown in Fig 2. We can see how, for all the models, regularization at the embedding level seems the preferable choice, even though regularizing the whole model does not impair performance equally in all the settings (e.g., STAtt).
Figure 7: Validation curves for different model families (5 runs, ±$\pm 1\sigma$). Plot names follow the convention model (embedding size) [dataset].
Similarly, Fig. 7 complements the results shown in Fig. 3, with the models that were not shown. We can notice a similar pattern in which dropout (red) can be problematic at smaller embedding sizes (top row). Nonetheless, this seems to be less significant for the STGNN in this specific scenario. In general, the additional figures confirm that dropout (red) and forgetting (blue) lead to different learning curves w.r.t. the un-regularized model (orange), while other regularizations have little impact on this aspect.
Figure 8: Validation curves for different model families (5 runs, ±$\pm 1\sigma$). Plot names follow the convention model (embedding size) [dataset].
Fig. 8 show results obtained by training models with the same setting as Sec. 5.2 on the PEMS-BAY dataset. We can see that, in this case, the validation curves barely plateau. In this scenario, dropout shows potentially problematic effects, similar to those observed for small embedding sizes.
Figure 9: Test performance degradation when adding zero-mean gaussian noise to the embeddings, with increasing variance (5 runs, ±$\pm 1\sigma$, METR-LA). Fixed hidden size and learning rate, $64$ and $0.00075$ respectively.
Figure 10: Test performance degradation on embeddings perturbation (5 runs, ±$\pm 1\sigma$, METR-LA).(Left) random shuffling, (Middle) replaced with their mean, and (Right) replaced by a draw from their sample normal. Fixed hidden size and learning rate, $64$ and $0.00075$ respectively.
Finally, Fig. 9 and Fig. 10, provide complementary results for Fig. 4. We can notice how the different regularizations rank similarly, in terms of global model robustness, across different architectures.